The market for a product has inverse demand and supply functions given by p = 290 - 2Qd and p = 10 + 1.5Qs

Find the market equilibrium quantity Q* and price P*.
Answer #1

Given inverse demand p=290-2Qd

Supply function p=10+1.5Qs

At equilibrium

Demand =supply

290-2Q=10+1.5Q

280=3.5Q

Q=80

Equilibrium quantity=80

   p=290-2(80)=290-160=$130

P=10+1.5(80)=10+120=$130

Equilibrium price=$130